OpenStudy (anonymous):

the cost of renting a jet ski is $40 per day plus $50 per hour of use. how many hours was a jet ski rented if the total cost was $390?

OpenStudy (amistre64):

we have a one time cost of 40 and a variable cost of 50 that depends on hours used... C(x) = 40 +50x = 390 0 = 40 + 50x -390 0 = 50x -350 x = 350/50 = 35/5 = 7